EXPERIMENTAL STUDY OF LINE PARAMETERS OF H_{2}S AROUND 1.57μm USING TELECOM DIODE LASER SPECTROSCOPY

A.-W. Liu, S.-M. Hu, Hefei National Laboratory for Physical Sciences at Microscale, University of Science and Technology of China, Hefei, 230026, China.; \underline{W.-D. Chen}, Laboratoire de Physicochimie de l’Atmosphère, Université du littoral Côte d’Opale, 189 AAν. Maurice Schumann, 59140 Dunkerque, France(chen@univ-littoral.fr).

Line intensities, self- and air-broadened linewidths of hydrogen sulfide (H_{2}S) have been determined around 1.57 μm utilizing Voigt line profiles. The measurements were performed by means of direct absorption spectroscopy technique using a continuous-wave tunable telecom-grade external cavity diode laser (ECDL) associated with a White type multipass cell. This spectral region corresponds to an interesting transparency window of the atmosphere. Transitions reaching the 2ν_{1}+ν_{2} and ν_{1}+ν_{2}+ν_{3} two upper vibrational states of H_{2}S bands were reported in our previous contribution^{a}. The line parameter analysis of 29 strong well-isolated lines from these two bands is extension of our knowledge.

aO. N. Ulenikov, A. W. Liu, E. S. Bekhtereva, O. V. Gromova, L. Y. Hao, S. M. Hu, J. Mol. Spectrosc. 234, 270, (2005)